Circular No. 6369

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 C. Kouveliotou, Universities Space Research Association; J. Greiner,

Max-Planck-Institut für Extraterrestrische Physik; J. van Paradijs, University of Alabama in Huntsville (UAH) and University of Amsterdam; G. J. Fishman, NASA/Marshall Space Flight Center; W. H. G. Lewin, R. Rutledge, and J. M. Kommers, Massachusetts Institute of Technology; and M. S. Briggs, UAH, report: “In a ROSAT HRI observation on Mar. 14 (exposure time 824 s), this source was detected at a mean countrate of 0.3 count/s, corresponding to roughly 1 mCrab in the band 0.1–2.4 keV. The best-fit x-ray position is α = 17^{h}44^{m}33^{s}.1, δ = −28^{o}44′29′′ (equinox 2000.0; estimated error ± 8′′). The variable radio source reported on IAUC 6307 and 6323 is 1′.5 from this x-ray position and thus not related to GRO J1744–28. Therefore, none of the candidate optical counterparts reported earlier are related to GRO J1744–28. Using a power-law model with photon index −1.22 and a column density N_{H} = 5.1 × 10^{22} cm^{−2}, the observed mean countrate corresponds to an unabsorbed luminosity of 3 × 10^{37} erg/s (D/10 kpc)^{2}. The error box of GRO J1744–28 was in the field-of-view of several ROSAT PSPC pointings in early 1992 March, during which the source was not detected. The deepest exposure gives a 3σ upper limit of 2 × 10^{−4} count/s (PSPC). Using the same spectral assumptions as above, this countrate limit corresponds to an unabsorbed luminosity limit of 4 × 10^{33} erg/s (D/10 kpc)^{2}.”

T. Augusteijn, European Southern Observatory (ESO); and van Paradijs, Kouveliotou, Greiner, Fishman, Rutledge, and Lewin further report: “We have taken a K-band image of the ROSAT error box of GRO J1744– 28 with the ESO New Technology Telescope on Mar. 28. A comparison of this image with the one taken on Feb. 8 by Blanco et al. (IAUC 6321) reveals the infrared counterpart of GRO J1744–28. The star is located at α = 17^{h}44^{m}33^{s}.1 ± 0^{s}.1, δ = −28^{o}44′19′′.5 ± 1′′.5 (equinox 2000.0), which is ∼ 9′′.5 due north of the center of the ROSAT error box reported by Kouveliotou et al. The star is detected at K = 15.7 ± 0.3 on Feb. 8, but it is undetected and at least 1 mag fainter on Mar. 28. The infrared brightness decrease indicates that GRO J1744–28 is a low-mass x-ray binary. Its K magnitude on Feb. 8 is consistent with that expected from an x-ray heated accretion disk in such a system with a 12-day orbital period and an xray luminosity of several times 10^{38} erg/s (Paradijs and McClintock 1994, A.Ap. 290, 133).”

1996 April 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green